there are100 people and 100 plates.among 100 people they are divided into three categories rich which take 2plates each,2 service class eat in 3 plates and 4 poor eat in 1 plate.calculate no of persons of each categories?
1
Expert's answer
2020-04-27T19:38:51-0400

We have x rich people, y service and z poor

And we have a system of 2 equations:

"x+y+z=100"

"2x+3\/2y+z\/4=100"

Which have this solutions

"y=60-7x\/5"

"z=40+2x\/5"

x,y,z should be integer and between 0 and 100. So we will have this possible options

1) "x=10, y=46, z=44"

2) "x=20, y=32, z=48"

3) "x=30, y=18, z=52"

4) "x=40, y=4, z=56"


Others won't be accepted because y should be divisible by 2, and z should be divisible by 4.